### v1

# Pressure suppresses the density wave order in kagome metal LuNb6Sn6

Abstract
Dancing tins pair up,
But compressing the framework
Thwarts the displacements.
The density waves that develop in kagome metals ScV6Sn6 and LuNb6Sn6 at low temperature appear to arise from under-filled atomic columns within a V-Sn or Nb-Sn scaffolding. Compressing this network with applied pressure in ScV6Sn6 suppressed the structural transition temperature by constraining atomic rattling and inhibiting the shifts that define the structural modulation. We predicted that the density wave transition in LuNb6Sn6 at 68 K would be suppressed by pressure as well. In this brief study we examine the pressure dependence of the density wave transition by remeasuring resistance vs temperature up to 2.26 GPa. We found the transition temperature is smoothly depressed and disappears around 1.9 GPa. This result not only addresses our prediction, but strengthens the rattling chains origin of structural instabilities in the HfFe6Ge6-type kagome metals.
I Introduction
Researchers have been enamored by the kagome lattice for the last decades due to the exciting behaviors that can arise in materials hosting this structural motif. In metallic systems, the connectivity of atoms on kagome sites generates electronic structures with desirable states like flat bands, van Hove singularities, and Dirac nodes.[1, 2]
Among metallic kagome compounds, structural modulations have emerged as a reoccurring theme. The CsV3Sb5-family,[3, 4] CsCr3Sb5,[5] V3Sb2,[6] La(Ru,Fe)3Si2,[7] Yb0.5Co3Ge3,[8] hexagonal-FeGe,[9, 10, 11] and ScV6Sn6 [12] all develop periodic structural distortions on cooling.
The density wave instability observed below 92 K in ScV6Sn6 has received significant attention due to some of its peculiar features. These include curious transport properties,[13, 14, 15, 16] competing density wave instabilities,[17, 18, 19, 20, 21, 22, 23] and subtle band structure reconstruction.[24, 25, 26, 27, 28, 29]
The structural modulation in ScV6Sn6 has been interpreted to not arise from the Peierls-like Fermi surface nesting common for charge density wave systems.[23, 30, 31, 17, 18, 19, 28] Instead, extra space in the HfFe6Ge6-type structure promotes rattling of Sc-Sn chains[32] and ultimately the development of static order driven by modulating bonding between Sn1 atoms.[33, 19] This explains why the density wave in ScV6Sn6 can be suppressed by pressure[34] and doping by larger rare earths.[32] In short, HfFe6Ge6-type compounds can have rattling instabilities whenever a small atom sits in a sufficiently large scaffolding.
We recently demonstrated this principle through our synthesis and examination of the extensive Nb6Sn6 family of compounds.[35] The rattling of atoms was enhanced with decreasing rare earth size until the a density wave was observed in LuNb6Sn6 at 68 K. This observation reinforced the rattling chain model developed for ScV6Sn6[32] and inspired an important prediction: We hypothesized that the density wave transition temperature of LuNb6Sn6 will be suppressed with pressure. We anticipate that rattling and Sn1-Sn1 bond modulation will be inhibited as the Nb-Sn framework is compressed as seen in ScV6Sn6.
In this paper we present an expedient test of this idea by measuring the temperature-dependent resistance of LuNb6Sn6 under pressure to 2.26 GPa. We clearly observe the transition signature in the resistance data march to lower temperatures as pressure is increased. The density wave in LuNb6Sn6 is suppressed by applied pressure, confirming our prediction. We further compare the pressure evolution of the density wave transition in ScV6Sn6 and Lu-doped ScV6Sn6 reported previously. Our work reaffirms the applicability of the rattling chain model in under-filled network compounds like the HfFe6Ge6-type materials.
II Methods
Single crystal growth of LuNb6Sn6 was performed using the self-flux technique.[35] Elemental reagents were combined in a ratio of 8:2:90 = Lu:Nb:Sn. We utilized Ames Lab Lu metals, Nb powder (Alfa, 99.9%), and Sn shot (Alfa, 99.9%). Reagents were placed into 5 mL Al2O3 (Canfield) crucibles fitted with a catch crucible/porous frit and sealed within fused quartz ampoules with approximately 0.6–0.7 atm of argon cover gas.[36] Samples were heated to 1150 °C and thermalized for 18 h before cooling to 900 °C at a rate of 1 °C/hr. Samples are subsequently centrifuged at 900 °C. Single crystals are small, well-faceted, gray-metallic, hexagonal plates and blocks. The samples are stable in air, water, and common solvents.
For these high-pressure measurements, a 4 mm ID hybrid cell was used with a BeCu outer wall and NiCrAl alloy for an insert. Copper wires and a fiber optic were fixed in place through the sample platform with Stycast 2850 FT epoxy. A hexagonal block-shaped crystal of LuNb6Sn6 was hand polished into a minuscule bar with a 80120 m cross section using sand paper. Contacts to the sample were made with 0.0005 in platinum wire in a conventional 4-contact arrangement held in place with Epotek H20E silver epoxy cured for 30 min at 130 °C. The pressure medium is Daphne 7575 oil which has been shown to remain hydrostatic to pressures above 3 GPa.[37] Transport measurements of the sample were made with a Lake Shore 372 resistance bridge with 3708 pre-amp/scanner using 1 mA and a 5 s averaging time. All data was collected with cooling curves taken at a rate of 0.4 K/min. Accurate temperatures were ensured by fixing a Cernox thermometer to the outside of the pressure cell and allowing for additional exchange gas in the sample space of the PPMS where the experiment took place. Resistance data was averaged over 15 measurements to reduce the noise in the derivative.
The fiber optics ends were cleaved cleanly to allow for a small chip of ruby to be glued to the end inside of the sample volume. Green laser light (532 nm) was transmitted to the ruby chip and the return fluorescence spectrum was determined by an Ocean Optics spectrometer. Ruby peaks are known to shift at a rate of 0.365 nm/GPa[38] and a room temperature and low temperature value for pressure were recorded for each curve. The reported pressures are from the region of interest (low temperatures) where the charge density wave was suppressed.
III Results
Figure 1a presents the normalized resistance vs temperature data obtained for a series of pressure. At the lowest pressure, resistivity drops sharply on cooling through 64 K. This correponds to the development of the density wave modulation observed at zero pressure at 68 K.[35] In ScV6Sn6 this drop has been tied to an increase in carrier mobility in the density wave phase.[12, 39]
The shape of the resistance curve evolves neatly with increasing pressure (lighter colors). The cliff-like drop at 0.1 GPa evolves into a cusp-shaped peak at 0.54 and 0.78 GPa. This further morphs into a broader step-up on cooling exemplified by the 1.26 GPa curve. The pressure evolution of this resistance feature mimics that observed in ScV6Sn6[34] and (Lu0.944Lu0.054)V6Sn6.[32] Importantly, the high-pressure jump-up feature is most dramatic in LuNb6Sn6.
There is a sharp drop in resistance between 2.8 and 3.7 K. A linear fit to the onset temperature vs pressure yields a slope of -0.410.01 K/GPa with an intercept of 3.740.02 K. This agrees exquisitely with the pressure dependence the superconducting critical temperature of tin metal ().[40] Despite careful polishing, there is a Sn inclusion in our sample.
The evolving nature of the density wave transition feature in this data set necessitates a systematic approach to extracting the transition temperatures. We plot the temperature derivative of resistance presented in Fig. 1b to assist us. We selected the maxima and minima in as characteristic temperatures as they correspond to the positive and negative slope extremes of . We identify these temperatures with estimated uncertainties with circles and squares in Fig. 1b.
At low pressures where a step-down feature is observed, the maximum of is best descriptor of this first-order transition temperature. In contrast, at higher pressure (like 1.26 GPa) the steepest negative slope of appears to be the best descriptor for a broadened first-order transition. At intermediate pressures, we could imagine the transition temperature at the cusp in which naturally lies between the maxima and minima of .
Importantly, we see no evidence of bulk superconductivity emerging as the density wave disappears. This was also true for ScV6Sn6 under pressure,[34] doped ScV6Sn6, and Lu-doped ScV6Sn6 under pressure.[25, 32] The first order nature of this transition in these compounds may provide a poor starting point for building a superconducting state.
Figure 2 presents the phase diagram of LuNb6Sn6 under pressure. The purple data clearly reveal that the transition temperature is smoothly suppressed with applied pressure and disappears by roughly 1.9 GPa. The solid line is included to emphasize the trend.
IV Discussion
We have confirmed our prediction that the density wave phase in LuNb6Sn6 is suppressed by pressure. This supports the rattling chains picture developed for ScV6Sn6[32] we extended to LuNb6Sn6.[35] Compressing the Nb-Sn scaffolding removes the extra space in the structure, inhibiting density wave formation.
Demonstrating broader applicability of the rattling chain model strengthens its value for materials design. Displacement fluctuations can be engineered by under-filling channels in compounds with rigid scaffolding networks. This could to enhance thermoelectric performance and induce density wave transitions. In addition, identifying the critical pressure required to prevent density wave formation will aid computational efforts exploring the criteria for structural instabilities in the HfFe6Ge6-type metals.
The pressure evolution of the transition temperatures in ScV6Sn6 and (Sc0.944Lu0.056)V6Sn6 are plotted fig. 2.[34] Curiously, although the Lu doped material has a significantly reduced transition temperature at zero pressure it shares a comparable critical pressure with the ScV6Sn6. LuNb6Sn6 has a notably lower critical pressure. This difference might be a result of the Nb-Sn framework being more compressible than the V-Sn. This should be addressable with DFT calculations or diffraction experiments under pressure.
Next we turn to a trend common to all three materials, the evolution of the transition signature shape with pressure. In most density wave phases the resistance is higher in the low temperature phase due to a loss of carrier when band reconstruct.[41] In ScV6Sn6, only the Sn1 p-bands show significant reconstruction accounting for a minor reduction in carrier concentration.[23, 24, 19, 33] It was proposed that the reduced resistance in the density wave phase is due to the strong scattering[12, 13, 39] by low energy phonons at temperatures just above the transition.[17, 18, 20] It is likely the same explanation lies behind the step down transition feature observed in LuNb6Sn6 at zero-pressure.
Why then does the transition step difference change sign on increasing pressure in ScV6Sn6 and LuNb6Sn6? First, this might represent a crossover in the relative impact of the change of carrier concentration and mobility as the density wave develops. Alternatively, there may be a change in the nature of the density wave order at higher pressures. For example, applying pressure might favor the competing 1/3 1/3 1/2 density wave[17, 18, 20] over the 1/3 1/3 1/3 observed at zero-pressure in both compounds.[12, 35] These ideas could be tested by low-temperature Raman spectroscopy, optical conductivity, or x-ray scattering under pressure.
V Conclusion
The density wave instabilities of ScV6Sn6 and LuNb6Sn6 are tied extra space provided by undersized atoms in a rigid V-Sn or Nb-Sn network. In the scandium compound, compressing this network with pressure inhibits the displacements that constitute the density wave, reducing the transition temperature. We have demonstrated that density wave order in LuNb6Sn6 is suppressed by pressure as we predicted. This observation emphasizes the value of the rattling chain model for density wave instabilities in filled network compounds.
VI Acknowledgments
Acknowledgements.
WRM acknowledges support from the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF9069 awarded to D.M.. Work by B.R.O. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. S.M. and D.M. acknowledge the support from AFOSR MURI (Novel Light-Matter Interactions in Topologically Non-Trivial Weyl Semimetal Structures and Systems), Grant No. FA9550-20-1-0322. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida.References
- Wang et al. [2013] W.-S. Wang, Z.-Z. Li, Y.-Y. Xiang, and Q.-H. Wang, Competing electronic orders on kagome lattices at van Hove filling, Physical Review B—Condensed Matter and Materials Physics 87, 115135 (2013).
- Kiesel et al. [2013] M. L. Kiesel, C. Platt, and R. Thomale, Unconventional Fermi surface instabilities in the kagome Hubbard model, Physical Review Letters 110, 126405 (2013).
- Ortiz et al. [2019] B. R. Ortiz, L. C. Gomes, J. R. Morey, M. Winiarski, M. Bordelon, J. S. Mangum, I. W. H. Oswald, J. A. Rodriguez-Rivera, J. R. Neilson, S. D. Wilson, E. Ertekin, T. M. McQueen, and E. S. Toberer, New kagome prototype materials: discovery of KV3Sb5,RbV3Sb5 , and CsV3Sb5, Physical Review Materials 3, 094407 (2019).
- Wilson and Ortiz [2024] S. D. Wilson and B. R. Ortiz, Sb5 kagome superconductors, Nat. Rev. Mater. , 9756 (2024).
- Li et al. [2024] Y. Li, Y. Liu, X. Du, S. Wu, W. Zhao, K. Zhai, Y. Hu, S. Zhang, H. Chen, J. Liu, Y. Yang, C. Peng, M. Hashimoto, D. Lu, Z. Liu, Y. Wang, Y. Chen, G. Cao, and L. Yang, Correlated electronic structure and incipient flat bands of the kagome superconductor CsCr3Sb5 10.48550/ARXIV.2406.03740 (2024), arXiv:2406.03740 [cond-mat.str-el] .
- Wang et al. [2022a] N. Wang, Y. Gu, M. A. McGuire, J. Yan, L. Shi, Q. Cui, K. Chen, Y. Wang, H. Zhang, H. Yang, X. Dong, K. Jiang, J. Hu, B. Wang, J. Sun, and J. Cheng, A density-wave-like transition in the polycrystalline V3Sb2 sample with bilayer kagome lattice, Chinese Physics B 31, 017106 (2022a).
- Plokhikh et al. [2024] I. Plokhikh, C. Mielke, H. Nakamura, V. Petricek, Y. Qin, V. Sazgari, J. Küspert, I. Biało, S. Shin, O. Ivashko, J. N. Graham, M. v. Zimmermann, M. Medarde, A. Amato, R. Khasanov, H. Luetkens, M. H. Fischer, M. Z. Hasan, J.-X. Yin, T. Neupert, J. Chang, G. Xu, S. Nakatsuji, E. Pomjakushina, D. J. Gawryluk, and Z. Guguchia, Discovery of charge order above room-temperature in the prototypical kagome superconductor La(Ru1-xFex)3Si2, Communications Physics 7, 10.1038/s42005-024-01673-y (2024).
- Wang et al. [2022b] Y. Wang, G. T. McCandless, X. Wang, K. Thanabalasingam, H. Wu, D. Bouwmeester, H. S. J. van der Zant, M. N. Ali, and J. Y. Chan, Electronic properties and phase transition in the kagome metal Yb0.5Co3Ge3, Chemistry of Materials 34, 7337 (2022b).
- Teng et al. [2023] X. Teng, J. S. Oh, H. Tan, L. Chen, J. Huang, B. Gao, J.-X. Yin, J.-H. Chu, M. Hashimoto, D. Lu, C. Jozwiak, A. Bostwick, E. Rotenberg, G. E. Granroth, B. Yan, R. J. Birgeneau, P. Dai, and M. Yi, Magnetism and charge density wave order in kagome FeGe, Nature Physics 19, 814 (2023).
- Teng et al. [2022] X. Teng, L. Chen, F. Ye, E. Rosenberg, Z. Liu, J.-X. Yin, Y.-X. Jiang, J. S. Oh, M. Z. Hasan, K. J. Neubauer, B. Gao, Y. Xie, M. Hashimoto, D. Lu, C. Jozwiak, A. Bostwick, E. Rotenberg, R. J. Birgeneau, J.-H. Chu, M. Yi, and P. Dai, Discovery of charge density wave in a kagome lattice antiferromagnet, Nature 609, 490 (2022).
- Chen et al. [2024] Z. Chen, X. Wu, S. Zhou, J. Zhang, R. Yin, Y. Li, M. Li, J. Gong, M. He, Y. Chai, X. Zhou, Y. Wang, A. Wang, Y.-J. Yan, and D.-L. Feng, Discovery of a long-ranged charge order with 1/4 Ge1-dimerization in an antiferromagnetic kagome metal, Nature Communications 15, 10.1038/s41467-024-50661-x (2024).
- Arachchige et al. [2022] H. W. S. Arachchige, W. R. Meier, M. Marshall, T. Matsuoka, R. Xue, M. A. McGuire, R. P. Hermann, H. Cao, and D. Mandrus, Charge density wave in kagome lattice intermetallic ScV6Sn6, Physical Review Letters 129, 216402 (2022).
- Mozaffari et al. [2024] S. Mozaffari, W. R. Meier, R. P. Madhogaria, N. Peshcherenko, S.-H. Kang, J. W. Villanova, H. W. S. Arachchige, G. Zheng, Y. Zhu, K.-W. Chen, K. Jenkins, D. Zhang, A. Chan, L. Li, M. Yoon, Y. Zhang, and D. G. Mandrus, Universal sublinear resistivity in vanadium kagome materials hosting charge density waves, Physical Review B 110, 035135 (2024).
- DeStefano et al. [2023] J. M. DeStefano, E. Rosenberg, O. Peek, Y. Lee, Z. Liu, Q. Jiang, L. Ke, and J.-H. Chu, Pseudogap behavior in charge density wave kagome material ScV6Sn6 revealed by magnetotransport measurements, npj Quantum Materials 8, 10.1038/s41535-023-00600-8 (2023).
- Yi et al. [2024] C. Yi, X. Feng, N. Mao, P. Yanda, S. Roychowdhury, Y. Zhang, C. Felser, and C. Shekhar, Quantum oscillations revealing topological band in kagome metal ScV6Sn6, Physical Review B 109, 035124 (2024).
- Kuo et al. [2024] C. N. Kuo, R. Y. Huang, W. S. Tian, C. K. Hong, Y. R. Ou, Y. K. Kuo, and C. S. Lue, Effects of lattice instability on the thermoelectric behavior of kagome metal ScV6Sn6, Applied Physics Letters 125, 10.1063/5.0231041 (2024).
- Cao et al. [2023] S. Cao, C. Xu, H. Fukui, T. Manjo, Y. Dong, M. Shi, Y. Liu, C. Cao, and Y. Song, Competing charge-density wave instabilities in the kagome metal ScV6Sn6, Nature Communications 14, 10.1038/s41467-023-43454-1 (2023).
- Korshunov et al. [2023] A. Korshunov, H. Hu, D. Subires, Y. Jiang, D. Călugăru, X. Feng, A. Rajapitamahuni, C. Yi, S. Roychowdhury, M. G. Vergniory, J. Strempfer, C. Shekhar, E. Vescovo, D. Chernyshov, A. H. Said, A. Bosak, C. Felser, B. A. Bernevig, and S. Blanco-Canosa, Softening of a flat phonon mode in the kagome ScV6Sn6, Nature Communications 14, 10.1038/s41467-023-42186-6 (2023).
- Hu et al. [2023a] H. Hu, Y. Jiang, D. Călugăru, X. Feng, D. Subires, M. G. Vergniory, C. Felser, S. Blanco-Canosa, and B. A. Bernevig, Kagome materials I: SG 191, ScV6Sn6. flat phonon soft modes and unconventional CDW formation: Microscopic and effective theory 10.48550/ARXIV.2305.15469 (2023a), arXiv:2305.15469 [cond-mat.str-el] .
- Pokharel et al. [2023] G. Pokharel, B. R. Ortiz, L. Kautzsch, S. J. Gomez Alvarado, K. Mallayya, G. Wu, E.-A. Kim, J. P. C. Ruff, S. Sarker, and S. D. Wilson, Frustrated charge order and cooperative distortions in ScV6Sn6, Physical Review Materials 7, 104201 (2023).
- Subedi [2024] A. Subedi, Order-by-disorder charge density wave condensation at =(1/3,1/3,1/3) in kagome metal ScV6Sn6, Physical Review Materials 8, 014006 (2024).
- Wang et al. [2024] K. Wang, S. Chen, S.-W. Kim, and B. Monserrat, Origin of competing charge density waves in kagome metal ScV6Sn6, Nat Commun 15, 10428 (2024) 15, 10.1038/s41467-024-54702-3 (2024), 2403.17058 [cond-mat.str-el] .
- Tan and Yan [2023] H. Tan and B. Yan, Abundant lattice instability in kagome metal ScV6Sn6, Physical Review Letters 130, 266402 (2023).
- Tuniz et al. [2023] M. Tuniz, A. Consiglio, D. Puntel, C. Bigi, S. Enzner, G. Pokharel, P. Orgiani, W. Bronsch, F. Parmigiani, V. Polewczyk, P. D. C. King, J. W. Wells, I. Zeljkovic, P. Carrara, G. Rossi, J. Fujii, I. Vobornik, S. D. Wilson, R. Thomale, T. Wehling, G. Sangiovanni, G. Panaccione, F. Cilento, D. Di Sante, and F. Mazzola, Dynamics and resilience of the unconventional charge density wave in ScV6Sn6 bilayer kagome metal, Communications Materials 4, 10.1038/s43246-023-00430-y (2023).
- Lee et al. [2024] S. Lee, C. Won, J. Kim, J. Yoo, S. Park, J. Denlinger, C. Jozwiak, A. Bostwick, E. Rotenberg, R. Comin, M. Kang, and J.-H. Park, Nature of charge density wave in kagome metal ScV6Sn6, npj Quantum Materials 9, 10.1038/s41535-024-00620-y (2024).
- Kim et al. [2023] D. W. Kim, S. Liu, C. Wang, H. W. Nam, G. Pokharel, S. D. Wilson, J.-H. Cho, and S. J. Moon, Infrared probe of the charge density wave gap in ScV6Sn6, Physical Review B 108, 205118 (2023).
- Cheng et al. [2024a] Z.-J. Cheng, S. Shao, B. Kim, T. A. Cochran, X. P. Yang, C. Yi, Y.-X. Jiang, J. Zhang, M. S. Hossain, S. Roychowdhury, T. Yilmaz, E. Vescovo, A. Fedorov, S. Chandra, C. Felser, G. Chang, and M. Z. Hasan, Untangle charge-order dependent bulk states from surface effects in a topological kagome metal ScV6Sn6 10.48550/ARXIV.2402.02341 (2024a), arXiv:2402.02341 [cond-mat.str-el] .
- Yang et al. [2024] Y.-C. Yang, S. Cho, T.-R. Li, X.-Q. Liu, Z.-T. Liu, Z.-C. Jiang, J.-Y. Ding, W. Xia, Z.-C. Tao, J.-Y. Liu, W.-C. Jing, Y. Huang, Y.-M. Shi, S. Huh, T. Kondo, Z. Sun, J.-S. Liu, M. Ye, Y.-L. Wang, Y.-F. Guo, and D.-W. Shen, Unveiling the charge density wave mechanism in vanadium-based bi-layered kagome metals 10.48550/ARXIV.2402.03765 (2024), arXiv:2402.03765 [cond-mat.mtrl-sci] .
- Kundu et al. [2024] A. K. Kundu, X. Huang, E. Seewald, E. Ritz, S. Pakhira, S. Zhang, D. Sun, S. Turkel, S. Shabani, T. Yilmaz, E. Vescovo, C. R. Dean, D. C. Johnston, T. Valla, T. Birol, D. N. Basov, R. M. Fernandes, and A. N. Pasupathy, Low-energy electronic structure in the unconventional charge-ordered state of ScV6Sn6, Nature Communications 15, 10.1038/s41467-024-48883-0 (2024).
- Cheng et al. [2024b] S. Cheng, Z. Ren, H. Li, J. S. Oh, H. Tan, G. Pokharel, J. M. DeStefano, E. Rosenberg, Y. Guo, Y. Zhang, Z. Yue, Y. Lee, S. Gorovikov, M. Zonno, M. Hashimoto, D. Lu, L. Ke, F. Mazzola, J. Kono, R. J. Birgeneau, J.-H. Chu, S. D. Wilson, Z. Wang, B. Yan, M. Yi, and I. Zeljkovic, Nanoscale visualization and spectral fingerprints of the charge order in ScV6Sn6 distinct from other kagome metals, npj Quantum Materials 9, 10.1038/s41535-024-00623-9 (2024b).
- Hu et al. [2024] Y. Hu, J. Ma, Y. Li, Y. Jiang, D. J. Gawryluk, T. Hu, J. Teyssier, V. Multian, Z. Yin, S. Xu, S. Shin, I. Plokhikh, X. Han, N. C. Plumb, Y. Liu, J.-X. Yin, Z. Guguchia, Y. Zhao, A. P. Schnyder, X. Wu, E. Pomjakushina, M. Z. Hasan, N. Wang, and M. Shi, Phonon promoted charge density wave in topological kagome metal ScV6Sn6, Nature Communications 15, 10.1038/s41467-024-45859-y (2024).
- Meier et al. [2023] W. R. Meier, R. P. Madhogaria, S. Mozaffari, M. Marshall, D. E. Graf, M. A. McGuire, H. W. S. Arachchige, C. L. Allen, J. Driver, H. Cao, and D. Mandrus, Tiny Sc allows the chains to rattle: Impact of Lu and Y doping on the charge-density wave in ScV6Sn6, Journal of the American Chemical Society 145, 20943 (2023).
- Liu et al. [2024] S. Liu, C. Wang, S. Yao, Y. Jia, Z. Zhang, and J.-H. Cho, Driving mechanism and dynamic fluctuations of charge density waves in the kagome metal ScV6Sn6, Physical Review B 109, l121103 (2024).
- Zhang et al. [2022] X. Zhang, J. Hou, W. Xia, Z. Xu, P. Yang, A. Wang, Z. Liu, J. Shen, H. Zhang, X. Dong, Y. Uwatoko, J. Sun, B. Wang, Y. Guo, and J. Cheng, Destabilization of the charge density wave and the absence of superconductivity in ScV6Sn6 under high pressures up to 11 GPa, Materials 15, 7372 (2022).
- Ortiz et al. [2024] B. R. Ortiz, W. R. Meier, G. Pokharel, J. Chamorro, F. Yang, S. Mozaffari, A. Thaler, S. J. G. Alvarado, H. Zhang, D. S. Parker, G. D. Samolyuk, J. A. M. Paddison, J. Yan, F. Ye, S. Sarker, S. D. Wilson, H. Miao, D. Mandrus, and M. A. McGuire, Stability frontiers in the kagome metals; the Nb6Sn6 (:Ce-Lu, Y) family and density-wave transition in LuNb6Sn6 10.48550/ARXIV.2411.10635 (2024), arXiv:2411.10635 [cond-mat.str-el] .
- Canfield et al. [2016] P. C. Canfield, T. Kong, U. S. Kaluarachchi, and N. H. Jo, Use of frit-disc crucibles for routine and exploratory solution growth of single crystalline samples, Philos. Mag. 96, 84 (2016).
- Staško et al. [2020] D. Staško, J. Prchal, M. Klicpera, S. Aoki, and K. Murata, Pressure media for high pressure experiments, Daphne oil 7000 series, High Pressure Research 40, 525 (2020).
- Piermarini et al. [1975] G. J. Piermarini, S. Block, J. D. Barnett, and R. A. Forman, Calibration of the pressure dependence of the ruby fluorescence line to 195 kbar, Journal of Applied Physics 46, 2774 (1975).
- Hu et al. [2023b] T. Hu, H. Pi, S. Xu, L. Yue, Q. Wu, Q. Liu, S. Zhang, R. Li, X. Zhou, J. Yuan, D. Wu, T. Dong, H. Weng, and N. Wang, Optical spectroscopy and band structure calculations of the structural phase transition in the vanadium-based kagome metal ScV6Sn6, Physical Review B 107, 165119 (2023b).
- Brandt and Ginzburg [1965] N. B. Brandt and N. I. Ginzburg, Effect of high pressure on the superconducting properties of metals, Soviet Physics Uspekhi 8, 202 (1965).
- Grüner [1994] G. Grüner, Density waves in solids (Perseus Publishing, Cambridge, Mass., 1994).